OpenStudy (anonymous):

What is 0.0314 written in scientific notation? No direct answers will medal and fan thanks you c: A. 3.14 × 10^–3 B. 3.14 × 10^–2 C. 3.14 × 10^2 D. 3.14 × 10^3

OpenStudy (anonymous):

In scientific notation, the first number has to be between 1 and 10. So 0.0314 = 3.14 /100 = 3.14 / 10^2 = ?

OpenStudy (anonymous):

Confused

OpenStudy (anonymous):

im not sure how else to explain it

OpenStudy (anonymous):

okay so we count the numbers behind 3?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

O that is what we do because i had someone else help me a sec ago and they were saying count the numbers behind it but i didnt know they had to be between 1-10

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I gave u a medal and im already a fan :3

OpenStudy (anonymous):

lol